OpenStudy (anonymous):

Find the values of b such that the function has the given maximum value. f(x) = −x2 + bx − 16; Maximum value: 65

OpenStudy (loser66):

That is a function of a downward parabola, so, the maximum of the function is its vertex. You have a formula to find it out, it says vertex has x = -b/2a = b/2 and replace that x into the function to have y = \(-\dfrac{b^2}{4} +b\dfrac{b}{2}-16 \) That is maximum value, so that y =65 So you have \(-\dfrac{b^2}{4} +b\dfrac{b}{2}-16=65 \) solve for b to get the answer